OpenStudy (anonymous):

Solve. 2= x-4/x-2 Possible Answers: x=-2 or x=-1 or x=0 or x=2

OpenStudy (mimi_x3):

2(x-2) = (x-4) 2x-4 = x-4 x = 0

OpenStudy (anonymous):

Thanks!!! :o)